The average American passenger car or light truck can drive about 20 miles on a gallon of gasoline: we say that its “fuel economy” is 20 miles per gallon. In the rest of the world, fuel economy would be expressed in kilometers per liter. Given that there are 1.6 kilometers in a mile and 3.8 liters in a gallon (approximately), express the fuel economy of the average American passenger vehicle in kilometers per liter


1
Expert's answer
2021-11-30T09:49:26-0500
"20\frac{mile}{gallon}=20\frac{1.6}{3.8}\frac{km}{L}=8.4\frac{km}{L}"